Determine if p(x) = 4x^2 - x is in the following span: W = span{x^2 + x, x^2 - 1, -x + 2}

(10 pts.) Determine if p(x) = 4x2 – x is in the following span: W = span {22 + x, x2 – 1, -«+2}.
Answer #1

p(x) = 4x² - x We span { x² + x, 22-1, - X+26 Now bra) 4x²-x = a (x²+x) + b1x²-1) + c(-x+2) 4x² - x = (a+b)x² + (a -c) x-bt a